OpenStudy (anonymous):

Can someone check my answers i got C What is the x-coordinate of the vertex for the parabola defined by the function y=-5x^2+9x+5? A.) -9/10 B.)-9 C.)9/10 D.)5/18

myininaya (myininaya):

y=-5(x^2-9x/5)+5+9 =-5(x-9/10)^2+14 9/10 is correct!